OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 A.x=0 B.x=2 C.There is no x value as there is no solution. D.x can be any value as there is an infinite number of solutions.

OpenStudy (solomonzelman):

divide the second equation by 2

OpenStudy (anonymous):

why would you divide

OpenStudy (solomonzelman):

please do

OpenStudy (anonymous):

So you divide every part by 2 or just a certain part

OpenStudy (solomonzelman):

when you divide the second equation by 2, you divide every term in the equation by 2.

OpenStudy (solomonzelman):

For example if I had 4x+6y=10 then I would get 2x+3y=5 after dividing by 2

OpenStudy (anonymous):

2x+y=1

OpenStudy (solomonzelman):

close, (-2) divided by 2 = -1

OpenStudy (anonymous):

So 2x+y=-1

OpenStudy (solomonzelman):

So, when we divide 4x+2y=-2 by 2, we get 2x+y=-1 (yes, correct)

OpenStudy (solomonzelman):

So, your new system is 2x+y=1 (your first equation) 2x+y=-1 (your second equation, after you divided it by 2)

OpenStudy (anonymous):

So then we add the equation together

OpenStudy (solomonzelman):

Now, tell me the following. According to the first equation what is \(\normalsize\color{royalblue}{ \rm 2x+y }\) is equal to \(\normalsize\color{royalblue}{ \rm 1 }\). According to the second equation what is \(\normalsize\color{royalblue}{ \rm 2x+y }\) equal to \(\normalsize\color{royalblue}{ \rm -1 }\).

OpenStudy (solomonzelman):

Regardless of what number \(\normalsize\color{royalblue}{ \rm 2x+y }\) is, can it be equal to 1 and -1 at the same time ?

OpenStudy (solomonzelman):

(if so, then 1=-1 )

OpenStudy (anonymous):

So it would be C.There is no x value as there is no solution.

OpenStudy (solomonzelman):

yes, there is no solution that satisfies the system (i.e. these equations are parallel lines)

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (solomonzelman):

anytime